# Fourier Neural Operators for Learning Dynamics in Quantum Spin Systems

### v1

@@ Abstract
Fourier Neural Operators (FNOs) excel on tasks using functional data, such as those originating from partial differential equations. Such characteristics render them an effective approach for simulating the time evolution of quantum wavefunctions, which is a computationally challenging, yet coveted task for understanding quantum systems. In this manuscript, we use FNOs to model the evolution of random quantum spin systems, so chosen due to their representative quantum dynamics and minimal symmetry. We explore two distinct FNO architectures and examine their performance for learning and predicting time evolution using both random and low-energy input states. Additionally, we apply FNOs to a compact set of Hamiltonian observables ($\sim\text{poly}(n)$) instead of the entire $2^{n}$ quantum wavefunction, which greatly reduces the size of our inputs and outputs and, consequently, the requisite dimensions of the resulting FNOs. Moreover, this Hamiltonian observable-based method demonstrates that FNOs can effectively distill information from high-dimensional spaces into lower-dimensional spaces. The extrapolation of Hamiltonian observables to times later than those used in training is of particular interest, as this stands to fundamentally increase the simulatability of quantum systems past both the coherence times of contemporary quantum architectures and the circuit-depths of tractable tensor networks.
@@
Figure 1: (a) Schematic representation of quantum spin system used to model its dynamics using FNOs, where $J_{x}$, $J_{y}$, and $J_{z}$ represent the coupling constants and $h$ denotes the external driving field. (b) Diagram illustrating the FNO framework, which consists of multiple FNO blocks performing spectral convolution over the input in the latent space. (c) We investigate two distinct architectures for learning dynamics in quantum spin systems. The first is the energy-domain architecture (see Section  I.2.1), where the input is the wavefunction at the initial time $t=0$, and the output is the wavefunction evolved to time $t=T$. (d) The second is the time-domain architecture (see Section  I.2.2), which takes the wavefunction evolved over the initial time interval $[0,\frac{3}{2}T]$ (discretized on a grid with width $\Delta t$) as input and produces an output wavefunction over the time interval $[T,\frac{5}{2}T]$.
@@
Our approach: We explore whether FNOs can effectively learn the dynamics of quantum wavefunctions in quantum systems, such as quantum spin systems [35]. Given the limitations of contemporary quantum simulators, FNOs present a potential tool for studying quantum systems, learning their dynamics and, most interestingly, extrapolating those dynamics to timescales that are not possible either experimentally (due to coherence times) or computationally (due to, e.g., the growing rank of tensor networks with increased circuit-depth or time-evolution). As quantum dynamics are dictated by the Schrödinger equation, which is a PDE, the use of FNOs for learning the time-evolution operator is well motivated. We employ temporal data from actual quantum computations to train FNOs, enabling them to capture system dynamics, extrapolate future time intervals, and generate predictions for analogous inputs. This methodology has the potential to address the limitations imposed by quantum decoherence, which impedes large-scale and prolonged simulations.
@@
In addition to using these two architectures on the full quantum wavefunction, we also deploy the time-domain architecture using a mere polynomial number of Hamiltonian observables as inputs and outputs. We are motivated to study this compact architecture because the full wavefunction analyses have utility only at scales for which classical simulations can provide a ground-truth calculation. Conversely, the more compact Hamiltonian observable training procedure can be applied to data from quantum devices that exceed the capabilities of classical simulations, representing a powerful future goal. This could enable the study of quantum systems that are too large to simulate classically on timescales that are too long to carry out experimentally. Another advantage of using Hamiltonian observables as inputs is that this can push the boundaries of tensor network simulation beyond what computationally-tractable bond dimensions, which increase linearly with the number of qubits but grow exponentially with circuit or unitary depth, can allow. This is because an FNO can be trained on data from shallower tensor networks and then be used to extrapolate to longer timescales, whereas simulating the same system directly with a deeper tensor network would result in computationally prohibitive bond dimensions. To demonstrate the potential of this goal, we extrapolate the dynamics of these Hamiltonian observables into timescales longer than those provided by the training data. This includes extending the Hamiltonian observable time horizon up to twice the duration of the training period, with a relative error of just $6.46\%$.
In the time-domain architecture, we achieve a substantial 6.71× speedup with FNOs compared to exact unitary evolution for inferring dynamics at later times for 8 qubit systems with only a minimal fidelity reduction of $0.04\%$. We note that this speedup is likely to become more substantial at larger system sizes, as both exact unitary integration and approximate integration techniques become more computationally intensive. Additionally, we demonstrate FNO’s capability for zero-shot super-resolution on 4 qubits by making predictions on a grid 10 times finer than that of the training interval. Due to its discretization-invariance, the FNO maintains exceptional accuracy, with error rates as low as 0.04% on the finer grid, while the U-Net shows a significantly increased error rate of 51.70% on the finer discretization. These results highlight FNO’s superior performance in both computational efficiency and accuracy, emphasizing its potential as a powerful tool for predicting quantum dynamics.
@@ I.1.1 Spin System Model
where $n$ represents the total number of qubits or atoms in the system. The Pauli matrix $\sigma_{i}^{a}$, where $a\in{x,y,z}$, is defined as $\sigma_{i}^{a}=I^{\otimes i-1}\otimes\sigma^{a}\otimes I^{\otimes n-i}$. Here, $I$ is the $2\times 2$ identity matrix and $\sigma^{a}$ denotes the corresponding $2\times 2$ Pauli matrix. The parameters $J_{x}$, $J_{y}$, and $J_{z}$ are the coupling constants for two-qubit spin interactions, while $h$ denotes the single-qubit driving field acting along the $z$-direction. We restrict interactions to nearest neighbors and apply periodic boundary conditions, such that $\sigma_{n+1}^{a}\equiv\sigma_{1}^{a}$. For our analysis, we use a Hamiltonian with randomly assigned coupling constants and a single-qubit driving field, where the values are uniformly distributed in the range from $-2$ to $2$. Additionally, we consider the quantum Ising Hamiltonian defined as
@@ I.1.2 Fourier Neural Operators
Neural Operators (NOs) are a class of machine learning models designed to learn mappings between infinite-dimensional function spaces, making them well-suited for a variety of applications, including ordinary differential equations (ODEs) and PDEs [38, 39, 40, 41]. A key advantage of NOs is their resolution-agnostic nature; they can be trained on data at one resolution and generalize to different resolutions without requiring retraining. Among Neural Operators, FNOs represent a specific implementation where spectral convolutions are utilized to capture the underlying patterns in the data [24].
@@
[…]
The lifting layer Lift maps the input data I with a given number of input channels to a latent function $\textbf{V}_{0}$ with a (typically) larger number of channels. Here, channels denote the dimensions of the co-domain of the functions $\textbf{V}_{\ell}$, representing distinct components or features within the latent space. The latent space is an intermediate representation where the data is abstracted into a higher-level form, capturing essential patterns and relationships. The FNO operates within this latent space by applying spectral convolutions.
@@
where $W_{\ell}$ is a learnable affine-linear map applied across the channels of $\textbf{V}_{\ell}$, and $\sigma$ is a non-linear activation function. The spectral convolution $K_{\ell}$ can be defined as follows,
@@
After processing through multiple FNO blocks $F_{\ell}$, the projection layer Proj maps the latent representation to the output data O, which may have one or more channels, depending on the specific architecture used as described in Sections I.2.1 and I.2.2.
@@ I.2 Learning Dynamics using Complete $2^{n}$ Wavefunction
In this approach, the full state space of a quantum wavefunction is used as input, with two distinct types of wavefunctions considered. The first type consists of complex-valued normalized wavefunctions for $n$ particles, randomly generated from a uniform distribution between $-1$ and $1$. For the second type, we uniformly distribute the wavefunction over low-energy states while setting the high-energy states to zero. This distinction is insightful, as in many physical applications, e.g., quantum chemistry [43], often only low-energy components of wavefunctions are occupied. In the remainder of this manuscript, we refer to the first input type as “random input” and the latter as “low-energy input”. We develop two distinct FNO architectures for processing wavefunction inputs: the energy-domain and time-domain architectures, which are described in detail below. Notably, because quantum wavefunctions are complex quantities, we use a complex version of the FNO [44].
@@ I.2.1 Energy-domain Architecture
The term $S_{0}$ corresponds to the basis states $S$ of a $2^{n}$ wavefunction, $\left[\left|\psi_{1}\right\rangle,\left|\psi_{2}\right\rangle,\cdots,\left|\psi_{2^{n}}\right\rangle\right]$, at the initial time $t=0$. To enhance the learning process, we incorporate a position embedding $\operatorname{Embed_{\mathrm{state}}}(k)=k/2^{n}$ into the input channel dimension, where $k$ is the index associated with each basis state. This approach helps the model better understand the relationship between the states and their corresponding energy levels. The output tensor is given by
@@
Thus the output, consists of wavefunction output evolved at $t=T$. In our experiments, we choose $T=\pi$ as it represents a significant portion of the quasi-periodicity of the unitary evolution generated by the Hamiltonian $H$. The Fourier transform on the basis states $S$ of the wavefunction is given by
$\frac{1}{2^{n}}\sum_{k=0}^{2^{n}-1}\psi_{k}e^{\frac{-ik\tau}{\hbar}},$ (7)
Here, $\psi_{k}$ is the basis state, $\tau$ can be understood as a time-like variable, and $\hbar$ is the reduced Planck constant. By truncating the transformed modes, we effectively filter out the components where energy values change rapidly, thus eliminating fast transitions. By concentrating on the slow-transition dynamics, the FNO highlights the most relevant dynamics of the quantum system. While this architecture provides a physical interpretation as a Fourier transform along the wavefunction’s energy space, it does not support further discretization of the time interval $[0,T]$, thereby limiting its utility in scenarios requiring more fine-grained prediction of the time-evolution. We will address this issue with our time-domain architecture in Section I.2.2.
(b) n Input Fidelity[$0,\pi$] Fidelity[$2\pi,10\pi$] 4 Random 0.9983 $\pm$ 0.0 0.9732 $\pm$ 2.3e-02 8 Random 0.9920 $\pm$ 0.0 0.9062 $\pm$ 5.8e-02 n Input Fidelity[$0,3\pi$] Fidelity[$4\pi,12\pi$] 4 Low Energy (VTI) 0.$9$$\pm$ 0.0 0.9992 $\pm$ 1.0e-03 8 Low Energy (VTI) 0.9996 $\pm$ 0.0 0.9430 $\pm$ 4.2e-02
Figure 2: (a) Prediction of temporal dynamics in an 8-qubit Heisenberg spin system using the energy-domain architecture in Section I.2.1. At time $T=\pi$, the predictions made by the FNO are on unseen data but still within the time range it was trained on. The extrapolated time refers to future time predictions beyond the range on which the FNO was trained. VTI (Various Time Interval) indicates training on multiple time intervals, i.e., $[0,T]$, $[T,2T]$, and $[2T,3T]$, instead of training only on the first interval. (b) The results show the mean fidelity for time intervals the FNO has seen during training and extrapolated time intervals for systems with 4 and 8 qubits.
@@
Experiments are conducted using $4$ and $8$ qubits with both random and low-energy input states. For low-energy states, training is conducted over various time intervals (VTI), such as $[0,T]$, $[T,2T]$, and $[2T,3T]$, rather than relying on a single time interval $[0,T]$. Utilizing multiple time intervals is crucial to generalize to future time intervals for low-energy states. Specifically, when training on only a single time step, the model fails to adequately capture the dynamics necessary for accurate future time predictions, despite showing strong performance during training. This is seen in Figure 2, where 8-qubit with low energy states has a mean fidelity of $0.7538$ for prediction of future time until $t=10T$ despite achieving a fidelity of $0.9998$ during training. This discrepancy underscores the importance of multi-step training for low-energy states, as these states initially occupy only low-energy configurations and gradually transition to higher-energy states over time. We test VTI when training models on random inputs. However, random inputs do not need to be trained across multiple intervals, as they already occupy a wide array of energy regions and can thus successfully extrapolate to later times with a fidelity greater than $94\%$. Figure 2 demonstrates that the model achieves high fidelity for both input types and can accurately predict future states while maintaining consistent fidelity across these predictions.

### v2

@@ Abstract
Fourier Neural Operators (FNOs) excel on tasks using functional data, such as those originating from partial differential equations. Such characteristics render them an effective approach for simulating the time evolution of quantum wavefunctions, which is a computationally challenging, yet coveted task for understanding quantum systems. In this manuscript, we use FNOs to model the evolution of random quantum spin systems, so chosen due to their representative quantum dynamics and minimal symmetry. We explore two distinct FNO architectures and examine their performance for learning and predicting time evolution using both random and low-energy input states. We find that standard neural networks in fixed dimensions, such as U-Net, exhibit limited ability to extrapolate beyond the training time interval, whereas FNOs reliably capture the underlying time-evolution operator and generalize effectively to unseen times. Additionally, we apply FNOs to a compact set of Hamiltonian observables ($\sim\text{poly}(n)$) instead of the entire $2^{n}$ quantum wavefunction, which greatly reduces the size of our inputs and outputs and, consequently, the requisite dimensions of the resulting FNOs. Moreover, this Hamiltonian observable-based method demonstrates that FNOs can effectively distill information from high-dimensional spaces into lower-dimensional spaces. Using this approach, we perform numerical experiments on a 20-qubit system and extrapolate Hamiltonian observables to twice the training time with a relative error of $5.8\%$. Notably, relative to numerical time-evolution methods, FNO achieves an inference speedup of approximately $10^{4}\times$ for 20-qubit systems, underscoring its computational efficiency at larger system sizes. The extrapolation of Hamiltonian observables to times later than those used in training is of particular interest, as this stands to fundamentally increase the simulatability of quantum systems past both the coherence times of contemporary quantum architectures and the circuit-depths of tractable tensor networks.
@@
Figure 1: (a) Schematic representation of quantum spin system used to model its dynamics using FNOs, where $J_{x}$, $J_{y}$, and $J_{z}$ represent the coupling constants and $h$ denotes the external driving field. (b) Diagram illustrating the FNO framework, which consists of multiple FNO blocks performing spectral convolution over the input in the latent space. (c) We investigate two distinct architectures for learning dynamics in quantum spin systems. The first is the energy-domain architecture (see Section I.2.1), where the input is the wavefunction at the initial time $t=0$, and the output is the wavefunction evolved to time $t=T$. (d) The second is the time-domain architecture (see Section I.2.2), which takes the wavefunction evolved over the initial time interval $0,\frac{3}{2}T]$ (discretized on a grid with width $\Delta t$) as input and produces an output wavefunction over the time interval $T,\frac{5}{2}T]$.
@@
Our approach: We explore whether FNOs can effectively learn the dynamics of quantum wavefunctions in quantum systems, such as quantum spin systems [34]. Given the limitations of contemporary quantum simulators, FNOs present a potential tool for studying quantum systems, learning their dynamics and, most interestingly, extrapolating those dynamics to timescales that are not possible either experimentally (due to coherence times) or computationally (due to, e.g., the growing rank of tensor networks with increased circuit-depth or time-evolution). As quantum dynamics are dictated by the Schrödinger equation, which is a PDE, the use of FNOs for learning the time-evolution operator is well motivated. We employ temporal data generated from high-fidelity classical simulations of quantum dynamics to train FNOs, enabling them to capture system evolution, extrapolate to future time intervals, and generalize for analogous inputs. This classical-data-driven approach ensures that the model learns accurate dynamics in a noise-free setting, while providing a foundation for future extensions that could use data directly from actual quantum computations to extrapolate dynamics beyond the coherence limit. This methodology has the potential to address the limitations imposed by quantum decoherence, which impedes large-scale and prolonged simulations.
@@
In addition to using these two architectures on the full quantum wavefunction, we also deploy the time-domain architecture using a mere polynomial number of Hamiltonian observables as inputs and outputs. We are motivated to study this compact architecture because the full wavefunction analyses have utility only at scales for which classical simulations can provide a ground-truth calculation. Conversely, the more compact Hamiltonian observable training procedure can be applied to data from quantum devices that exceed the capabilities of classical simulations, representing a powerful future goal. This could enable the study of quantum systems that are too large to simulate classically on timescales that are too long to carry out experimentally. Another advantage of using Hamiltonian observables as inputs is that this can push the boundaries of tensor network simulation beyond what computationally-tractable bond dimensions, which increase linearly with the number of qubits but grow exponentially with circuit or unitary depth, can allow. This is because an FNO can be trained on data from shallower tensor networks and then be used to extrapolate to longer timescales, whereas simulating the same system directly with a deeper tensor network would result in computationally prohibitive bond dimensions. To demonstrate the potential of this goal, we extrapolate the dynamics of these Hamiltonian observables into timescales longer than those provided by the training data. This includes extending the Hamiltonian observable time horizon for a $20$-qubit system up to twice the duration of the training period, with a relative error of just $5.8\%$.
In the time-domain architecture with wavefunction inputs, we achieve a substantial 6.71x speedup with FNOs compared to exact unitary evolution for inferring dynamics at later times for 8 qubit systems with only a minimal fidelity reduction of $0.04\%$. For Hamiltonian-observable inputs, the FNO demonstrates an even more remarkable inference speedup on the order of $10^{4}$x for 20-qubit systems, highlighting its efficiency for larger-scale quantum simulations. We note that this speedup is likely to become more substantial at larger system sizes, as both exact unitary integration and approximate integration techniques become more computationally intensive. Additionally, we demonstrate FNO’s capability for zero-shot super-resolution on 4 qubits by making predictions on a grid 10 times finer than that of the training interval. Due to its discretization-invariance, the FNO maintains exceptional accuracy, achieving error rates as low as 0.04% on the finer grid. For comparison, we include a U-Net baseline, a widely used convolutional encoder–decoder architecture with skip connections [46]. Due to its fixed-size kernels, U-Net learns patterns at a fixed resolution and does not learn the underlying operator [27]. When evaluated on the same finer discretization, the U-Net exhibits a substantially higher error rate of 51.70%. This contrast highlights that while U-Net can reproduce patterns at fixed resolution, it does not learn the underlying operator, emphasizing the advantage of FNOs for capturing system dynamics and enabling accurate extrapolation. As shown in Figure 3, the FNO markedly outperforms the U-Net in the extrapolation regime. The U-Net performs poorly, particularly on low-energy wavefunction inputs, with a mean fidelity of $0.1501$, whereas the FNO captures the underlying structure and extrapolates with high accuracy, achieving a mean fidelity of $0.8893$. Similarly, a direct comparison with classical trajectory-based methods such as Dynamical Mode Decomposition (DMD) is not feasible. DMD assumes knowledge of the system’s initial state and operates on snapshot pairs from a single trajectory, learning a linear operator that describes the evolution of that specific state [48]. In contrast, in our setting, the initial state is unknown, so DMD cannot be directly applied. The FNO, on the other hand, is trained on a diverse set of randomly generated quantum wavefunctions and learns a functional mapping of the time evolution across an entire class of states. This allows FNOs not only to reproduce local features but also to capture the underlying operator governing quantum dynamics, a capability that neither purely pattern-based ML models nor single-trajectory linear methods can provide. These results highlight FNO’s superior performance in both computational efficiency and accuracy, emphasizing its potential as a powerful tool for predicting quantum dynamics.
@@ I.1.1 Spin System Model
where $n$ represents the total number of qubits or atoms in the system. The Pauli matrix $\sigma_{i}^{a}$, where $a\in{x,y,z}$, is defined as $\sigma_{i}^{a}=I^{\otimes i-1}\otimes\sigma^{a}\otimes I^{\otimes n-i}$. Here, $I$ is the $2\times 2$ identity matrix and $\sigma^{a}$ denotes the corresponding $2\times 2$ Pauli matrix. The parameters $J_{x}$, $J_{y}$, and $J_{z}$ are the coupling constants for two-qubit spin interactions, while $h$ denotes the single-qubit driving field acting along the $z$-direction. We restrict interactions to nearest neighbors and apply periodic boundary conditions, such that $\sigma_{n+1}^{a}\equiv\sigma_{1}^{a}$. For our analysis, we use a Hamiltonian with randomly assigned coupling constants and a single-qubit driving field, where the values are uniformly distributed in the range from 2-2 to $2$. Additionally, we consider the quantum Ising Hamiltonian defined as
@@ I.1.2 Fourier Neural Operators
Neural Operators (NOs) are a class of machine learning models designed to learn mappings between infinite-dimensional function spaces, making them well-suited for a variety of applications, including ordinary differential equations (ODEs) and PDEs [2, 36, 17, 3]. A key advantage of NOs is their resolution-agnostic nature; they can be trained on data at one resolution and generalize to different resolutions without requiring retraining. Among Neural Operators, FNOs represent a specific implementation where spectral convolutions are utilized to capture the underlying patterns in the data [24, 8]. Although the input and output are represented on discrete grids during training, the Fourier Neural Operator (FNO) learns a mapping between functions rather than between fixed-length vectors. The lifting and projection layers act pointwise with shared parameters, while the Fourier layers capture global structure in the Fourier space. Together, these components define a continuous, grid-independent operator through its Fourier representation. As a result, the trained operator can be evaluated on discretizations finer than those used in training, enabling zero-shot super-resolution without additional training. This mechanism is closely analogous to spectral interpolation on regular grids, where samples of a periodic, band-limited function defined on an $N$-point uniform grid are transformed to Fourier space, symmetrically zero-padded to a larger spectral resolution $M>N$, and then transformed back via an inverse Fourier transform to obtain the function evaluated on a $M$-point finer grid. This procedure preserves the physically meaningful low-frequency modes without introducing any artificial high-frequency modes, resulting in the evaluation of the same underlying function on a finer grid [8]. This mechanism underlies the FNO’s discretization invariance, with the operator being grid-independent and the number of training points affecting only the numerical representation, not the learned mapping.
@@
$\textbf{I}\overset{\text{Lift}}{\longrightarrow}\textbf{V}_{0}\overset{F_{1}}{\longrightarrow}\textbf{V}_{1}\overset{F_{2}}{\longrightarrow}\ldots\overset{F_{L-1}}{\longrightarrow}\textbf{V}_{L-1}\overset{F_{L}}{\longrightarrow}\textbf{V}_{L}\overset{\text{Proj}}{\longrightarrow}\textbf{O}.$
The lifting layer Lift is a local transformation parameterized by a shallow fully connected neural network,
$\mathrm{Lift}:\mathbb{R}^{d_{i}}\rightarrow\mathbb{R}^{d_{v}},$
which embeds the input data $\mathbf{I}\in\mathbb{R}^{d_{i}}$ into a higher-dimensional latent channel space. This produces the initial latent representation
$\mathbf{V}_{0}=\mathrm{Lift}(\mathbf{I}),\qquad\mathbf{V}_{0}\in\mathbb{R}^{d_{v}},$
where typically $d_{v}>d_{i}$. Here, channels denote the dimensions of the co-domain of the functions $\textbf{V}_{\ell}$, representing distinct components or features within the latent space. The latent space is an intermediate representation where the data is abstracted into a higher-level form, capturing essential patterns and relationships. The FNO operates within this latent space by applying spectral convolutions [24, 8].
@@
where $W_{\ell}$ is a learnable affine-linear map applied across the channels of $\textbf{V}_{\ell}$, and $\sigma$ is a non-linear activation function. For all FNO experiments in this work, we employ the Gaussian Error Linear Unit (GELU) as the activation function. We choose GELU because it has been shown to perform well in smooth operator learning tasks and is more flexible than ReLU, allowing small negative outputs and being conducive to more stable training [8]. The spectral convolution $K_{\ell}$ can be defined as follows,
@@
After processing through multiple FNO blocks $F_{\ell}$, the projection layer Proj maps the latent representation back to the output function space.
$\mathrm{Proj}:\mathbb{R}^{d_{v}}\rightarrow\mathbb{R}^{d_{o}},$
which yields the predicted output
$\mathbf{O}=\text{Proj}(\mathbf{V}_{L}),\qquad\mathbf{O}\in\mathbb{R}^{d_{o}}.$
The output data O may have one or more channels, depending on the specific architecture used as described in Sections I.2.1 and I.2.2.
@@ I.2 Learning Dynamics using Complete $2^{n}$ Wavefunction
In this approach, the full state space of a quantum wavefunction is used as input, with two distinct types of wavefunctions considered. The first type consists of complex-valued normalized wavefunctions for $n$ particles. Each amplitude is generated by independently sampling its real and imaginary parts from a uniform distribution on the interval $[0,1)$ using torch.rand. The resulting wavefunction is then normalized to have unit $L^{2}$ norm. For the second type, we uniformly distribute the wavefunction over low-energy states while setting the high-energy states to zero. The exact method is described in detail in Section III. This distinction is insightful, as in many physical applications, e.g., quantum chemistry [21], often only low-energy components of wavefunctions are occupied. In the remainder of this manuscript, we refer to the first input type as “random input” and the latter as “low-energy input”. We develop two distinct FNO architectures for processing wavefunction inputs: the energy-domain and time-domain architectures, which are described in detail below. Notably, because quantum wavefunctions are complex quantities, we use a complex version of the FNO [43].
@@ I.2.1 Energy-domain Architecture
Here, $S_{0}$ is the vector of complex amplitudes of the $2^{n}$ wavefunction in the computational basis at initial time $t=0$, $S_{0}=\left[c_{1},c_{2},\dots,c_{2^{n}}\right]$, with $\sum_{i=1}^{2^{n}}|c_{i}|^{2}=1.$ To enhance the learning process, we incorporate a position embedding $\operatorname{Embed_{\mathrm{state}}}(k)=k/2^{n}$ into the input channel dimension, where $k$ is the index associated with each basis state. In this setting, the ordering of the basis states corresponds to increasing energy levels. By providing this positional embedding, the FNO can more effectively capture the hierarchical structure of the wavefunction, distinguishing between low- and high-energy components. This is particularly relevant because the FNO operates in Fourier space, where high-frequency modes are truncated. The position embedding ensures that these truncated modes predominantly correspond to high-energy components that contribute minimally to the system’s dynamics, thereby enhancing both the learning efficiency and extrapolation accuracy. The output tensor is given by
@@
Thus the output, consists of wavefunction output evolved at $t=T$. In our experiments, we choose $T=\pi$ as it represents a significant portion of the quasi-periodicity of the unitary evolution generated by the Hamiltonian $H$. The Fourier transform on the amplitudes $c_{k}$ of the wavefunction in the computational basis is given by
$\frac{1}{2^{n}}\sum_{k=0}^{2^{n}-1}c_{k}e^{\frac{-ik\tau}{\hbar}},$ (7)
where $c_{k}$ is the complex amplitude of the $k$-th computational basis state, $\tau$ can be understood as a time-like variable, and $\hbar$ is the reduced Planck constant. By truncating the transformed modes, we effectively filter out the components where energy values change rapidly, thus eliminating fast transitions. By concentrating on the slow-transition dynamics, the FNO highlights the most relevant dynamics of the quantum system. While this architecture provides a physical interpretation as a Fourier transform along the wavefunction’s energy space, it does not support further discretization of the time interval $[0,T]$, thereby limiting its utility in scenarios requiring more fine-grained prediction of the time-evolution. We will address this issue with our time-domain architecture in Section I.2.2.
Random initial states
n $F(\pi)^{\mathrm{train}}$ $F(2\pi)^{\mathrm{ext}}$ $F(6\pi)^{\mathrm{ext}}$ $F(10\pi)^{\mathrm{ext}}$ 4 $0.9997\pm 0.0001$ $0.9991\pm 0.0003$ $0.9949\pm 0.0026$ $0.9894\pm 0.0060$ 8 $0.9911\pm 0.0009$ $0.9762\pm 0.0023$ $0.9010\pm 0.0074$ $0.8101\pm 0.0129$
Low-energy states (VTI)
n $F(3\pi)^{\mathrm{train}}$ $F(4\pi)^{\mathrm{ext}}$ $F(8\pi)^{\mathrm{ext}}$ $F(12\pi)^{\mathrm{ext}}$ 4 $99\pm 0.0001 0.9999±0$$99\pm 0.0001$ $0.9994\pm 0.0005$ $0.9984\pm 0.0014$ 8 $0.9984\pm 0.0002$ $0.9806\pm 0.0047$ $0.9202\pm 0.0125$ $0.8836\pm 0.0184$
Figure 2: (a) Prediction of temporal dynamics in an 8-qubit Heisenberg spin system using the energy-domain architecture in Section I.2.1. At time $T=\pi$, the predictions made by the FNO are on unseen data but still within the time range it was trained on. The extrapolated time refers to future time predictions beyond the range on which the FNO was trained. VTI (Various Time Interval) indicates training on multiple time intervals, i.e., $[0,T]$, $[T,2T]$, and $[2T,3T]$, instead of training only on the first interval. (b) The results show the fidelity at specific time steps for 4 and 8 qubits, including times within the training range as well as extrapolated future times beyond the training interval. Superscripts indicate whether the particular time step is within the training range ($\mathrm{train}$) or an extrapolated time ($\mathrm{ext}$).
@@
Experiments are conducted using $4$ and $8$ qubits with both random and low-energy input states. For low-energy states, training is conducted over various time intervals (VTI), such as $[0,T]$, $[T,2T]$, and $[2T,3T]$, rather than relying on a single time interval $[0,T]$. Utilizing multiple time intervals is crucial to generalize to future time intervals for low-energy states. Specifically, when training on only a single time step, the model fails to adequately capture the dynamics necessary for accurate future time predictions, despite showing strong performance during training. This is seen in Figure 2, where 8-qubit with low energy states has a mean fidelity of $0.7538$ for prediction of future time until $t=10T$ despite achieving a fidelity of $0.9998$ during training. This discrepancy underscores the importance of multi-step training for low-energy states, as these states initially occupy only low-energy configurations and gradually transition to higher-energy states over time. When VTI is applied, low-energy inputs achieve substantially higher mean fidelities at extrapolated times, reaching $0.9992$ for 4-qubit and $0.9430$ for 8-qubit states. We also test VTI when training models on random inputs. Random inputs do not strictly require multiple intervals, as they already occupy a wide array of energy regions and can successfully extrapolate to later times. For example, 4-qubit random inputs without VTI reach a mean fidelity of about $97\%$, while 8-qubit random inputs achieve around $91\%$. Nonetheless, applying VTI even for random inputs further improves performance, achieving mean fidelities as high as $99.99\%$ for 8 qubits, highlighting that multi-interval training can still be beneficial.
FNO performs better on random wavefunctions as these states already span a wide range of energy modes and configurations. Each random state effectively provides the model with information across the entire Hilbert space, allowing it to learn the dynamics for all relevant energy regions in a single step. In contrast, low-energy states are initially localized to only a small subset of modes, and accurately predicting their future evolution requires the model to capture how population spreads to higher-energy modes over time. This highlights why VTI is necessary for accurately extrapolating low-energy states, whereas random states achieve satisfactory results even when the model is trained on a single time step. Figure 2 demonstrates that the model achieves high fidelity for both input types and can accurately predict future states while maintaining consistent fidelity across these predictions.